# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ Abstract
We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time fixed effects in both the intercept and slope. These models are subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at zero, and that test statistics do not follow their standard asymptotic distributions as in the absence of the fixed effects. To address the problem, we develop an analytical bias correction procedure to construct a bias-corrected likelihood. The resulting estimator follows an asymptotic normal distribution with mean zero. Moreover, likelihood-based tests statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-squared distribution under the null hypothesis. Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
@@ 1 Introduction
Panel data is common in empirical research. Panel data models with two-way fixed effects are widely used to control for unobserved heterogeneity that may correlate with covariates. Individual effects capture time-invariant, individual-specific characteristics, while time effects account for shifts that are common across individuals but vary over time. Traditionally, these models address level heterogeneity by incorporating fixed effects only in the intercept, adjusting for baseline outcome differences across individuals and time periods. However, many empirical applications reveal substantial variation in how outcomes respond to covariates (Browning and Carro, 2007). When such response heterogeneity are present and correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
@@
To illustrate, consider the classic problem of estimating the impact of children on a mother’s labor force participation. Research indicates that mothers with more children may differ systematically from those with fewer or none (Blau and Robins 1988; Angrist and Evans 1998; Kleven et al. 2019). First, mothers who choose to have more children might be inherently less inclined to work. Second, mothers with more children may be those whose labor supply is less sensitive to family size, perhaps due to better financial resources or access to childcare. In a panel dataset tracking households over time, we can control for the first type of level heterogeneity by including fixed effects in the intercept. However, to address the second type of response heterogeneity, we must include fixed effects in the slope. Given that labor force participation is a binary outcome and typically exhibits serial correlation, this example underscores the need for a dynamic nonlinear panel data model with fixed effects in both the intercept and slope.
@@
In this paper, we introduce a general class of dynamic nonlinear heterogeneous parameter models (DN-HP) that incorporate individual and time fixed effects in both the intercept and slope. This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel data models. DN-HP models are generally subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at 00, and that the test statistics do not follow their standard asymptotic distributions (of $\chi^{2}$, for many test statistics). To address the incidental parameter problem, we propose an analytical bias correction approach. This method restores the standard asymptotic distributions for the point estimator and the test statistics.
@@
Notes: Model. $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})X_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of $X_{it}$, and the true values $\rho_{0}=\beta_{0}=0.5$. $\widehat{\beta}$ and $\widehat{\rho}$ are the uncorrected estimators while $\widehat{\beta}_{L}$ and $\widehat{\rho}_{L}$ are the corrected estimators. $1000$ replications. The detailed data-generating process is given in Section 4.
@@
As a preview, Figure 1 presents simulated boxplots of the estimation bias in a DN-HP logit model, comparing estimators from our bias correction procedure to the uncorrected ones. The results reveal that our bias correction procedure significantly reduces the estimation bias while maintaining comparable mean square errors.
@@
Literature Review. The estimation and inference of fixed effects models in the presence of the incidental parameter problem have been extensively studied. Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, addressing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models. In recent decades, the availability of long panel datasets has motivated a large-$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow proportionally to infinity. Under this rectangular asymptotic, estimators remain consistent but typically exhibit a non-negligible bias of order $O(1/T)$ due to the presence of individual effects in the intercept. To enable valid inference, researchers have developed various bias correction methods for a broad range of models in a maximum likelihood framework. For instance, Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” bias correction methods to eliminate the bias from the maximum likelihood estimator. Two alternative approaches have also been explored: “score-based” and “likelihood-based” methods. Woutersen (2002) and Li et al. (2003) modify the profiled score to remove the bias, while Bester and Hansen (2009) and Arellano and Hahn (2016) identify and correct the bias directly for the likelihood. Most of these methods are analytical, estimating the bias using explicit formulas. Alternatively, numerical procedures are available, estimating the bias through methods such as the Jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023), and integrated likelihoods (Arellano and Bonhomme, 2009).
@@
For two-way fixed effects models, incorporating time effects in the intercept generates an additional bias of $O(1/N)$. Moon and Weidner (2017) propose a parameter-based analytical correction for dynamic linear models with two-way interactive effects. For nonlinear models, Fernández-Val and Weidner (2016) develop parameter-based bias correction techniques with additive fixed effects, while Chen et al. (2021) focus on interactive fixed effects in static models. Alternatively, Jochmans and Otsu (2019) propose a likelihood-based analytical bias correction for static models, allowing fixed effects to enter the model in an arbitrary but known form. Other model-specific solutions include Bai (2009) and Charbonneau (2017), among others. For an extensive review of the incidental parameter problem for models with two-way fixed effects in the intercept, see Fernández-Val and Weidner (2018).
To our knowledge, our paper represents one of the first efforts to extend intercept-only two-way fixed effects models into multi-dimensional two-way fixed effects models, accounting for both level and response heterogeneities. Two prior studies in this line of research are Keane and Neal (2020) and Lu and Su (2023). Both examine linear heterogeneous parameter models with additive effects in both the intercept and slope. Keane and Neal (2020) propose an iterative procedure, which they refer to as the “mean observation OLS,” to eliminate the bias, while Lu and Su (2023) rely on a parameter-based Jackknife approach and provide a uniform inference theory for structural parameters. Our study significantly expands the model framework and introduces a likelihood-based analytical bias correction technique to ensure valid estimation and inference. Neither our DN-HP models nor our bias correction procedure have been discussed in the current literature.
@@
This paper. We employ the maximum likelihood framework and focus on multi-dimensional additive two-way fixed effects, whose number grows with the sample size. We consider an arbitrary DN-HP model whose log-likelihood function is specified up to the unknown parameters. We construct a modified, or corrected, log-likelihood function by adding two bias correction terms to the original one. These two bias terms are analytically derived by combining i) the Taylor expansion of the original log-likelihood function, in terms of the fixed effects, and ii) an asymptotic expansion of the fixed effects themselves. We show that the corrected log-likelihood function does not suffer from the incidental parameter problem, under appropriate regularity conditions.
@@
One key benefit of our procedure is that we achieve bias corrections for the point estimators and the test statistics by modifying a single object, the log-likelihood function. We rigorously show that the estimators obtained by maximizing the corrected log-likelihood function retains a limiting normal distribution with zero mean, consistent with the asymptotic theory of the classical maximum likelihood estimators (MLE) in the absence of incidental parameters. Beyond the estimators, we also show that the likelihood ratio (LR), Lagrange multiplier (LM), and Wald statistics, derived from the same corrected log-likelihood function, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
@@
We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to both the original likelihood and the Jackknife, which is a potential bias correction device [^1] It is worth mentioning that, to our knowledge, there currently has been no study on using the jackknife under the DN-HP models.. We show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors, and restores the test sizes of the LR test to its nominal level. We also find that our procedure outperforms the Jackknife. In the empirical application, we apply the DN-HP logit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households. Our analysis reveals that this impact varies significantly across individuals and over time. Neglecting such variation can lead to biased estimates and markedly different conclusions, highlighting the need to address both level and response heterogeneities, as well as the value of our framework in empirical research.
@@
The remainder of the paper is organized as follows. Section 2 explains our settings, the identification restriction, and the incidental parameter problem of DN-HP models. Section 3 describes our bias correction procedure and provides relevant statistical properties, including those we claim above. Section 4 presents simulation studies to demonstrate the performance of method. Section 5 applies our method to study single mothers’ labor force participation. Finally in Section 6, we leave some closing remarks. Proofs, technical details, and additional results are provided in the appendix.
@@ 2.1 Models and Estimations
$Y_{it}|X_{i1},\ldots,X_{it},\phi^{0}\sim f(y|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$ (2.1)

### v5

@@ Abstract
We introduce a generic class of dynamic nonlinear heterogeneous parameter models that incorporate individual and time fixed effects in both the intercept and slope. These models are subject to the incidental parameter problem, in that the limiting distribution of the point estimator is not centered at zero, and that test statistics do not follow their standard asymptotic distributions as in the absence of the fixed effects. To address the problem, we develop an analytical bias correction procedure to construct a bias-corrected likelihood. The resulting estimator follows an asymptotic normal distribution with mean zero. Moreover, likelihood-based test statistics—including likelihood-ratio, Lagrange-multiplier, and Wald tests—follow the limiting chi-squared distribution under the null hypothesis. Simulations demonstrate the effectiveness of the proposed correction method, and an empirical application on the labor force participation of single mothers underscores its practical importance.
@@ 1 Introduction
Panel data are common in empirical research. Panel data models with two-way fixed effects are widely used to control for unobserved heterogeneity that may correlate with covariates. Individual effects capture time-invariant, individual-specific characteristics, while time effects account for shifts that are common across individuals but vary over time. Traditionally, these models address level heterogeneity by incorporating fixed effects only in the intercept, adjusting for baseline outcome differences across individuals and time periods. However, both economic theory and empirical evidence indicate substantial variation in how outcomes respond to covariates across individuals and over time (Browning and Carro, 2007). Cross-sectional response heterogeneity reflects individual-specific factors such as preferences, productivity, and access to resources (Heckman, 2001), whereas time-varying response heterogeneity captures changes in behavior driven by evolving circumstances—including policy reforms, business cycles, and technological progress (Okui and Wang, 2021). When such response heterogeneities are correlated with covariates, failing to account for them can lead to biased estimates and invalid inference (Su and Chen, 2013).
@@
To illustrate, consider the classic problem of estimating the impact of children on a mother’s labor force participation. Research indicates that mothers with more children may differ systematically from those with fewer or none (Blau and Robins, 1988; Angrist and Evans, 1998; Kleven et al., 2019). First, mothers who choose to have more children might be inherently less inclined to work. Second, mothers with more children may be those whose labor supply is less sensitive to family size, perhaps due to better financial resources or childcare access. Moreover, over time, declining fertility and rising female labor force participation suggest a shifting baseline, while advances in home technology and welfare reforms may lessen the burden of additional children, enhancing labor supply responsiveness. In a panel dataset tracking households over time, we can control for the first type of heterogeneity—level differences in baseline participation—by including fixed effects in the intercept. However, to address the second type of heterogeneity—variations in response to children across individuals and over time—we must include fixed effects in the slope. Given that labor force participation is a binary outcome and typically exhibits serial correlation, this example underscores the need for a dynamic nonlinear panel data model with fixed effects in both the intercept and slope.
@@
In this paper, we introduce a general class of dynamic nonlinear heterogeneous-parameter (DN-HP) models that incorporate individual and time fixed effects in both the intercept and the slope. This flexible framework accommodates multi-dimensional heterogeneities and encompasses many commonly used panel-data models, including both static and dynamic specifications for linear and limited-dependent-variable outcomes. It nests individual-specific slope models that capture heterogeneous responses across individuals, as well as time-varying-coefficient models that allow response parameters to evolve over time due to shifts in the underlying economic or structural environment. DN-HP models are generally subject to the incidental-parameter problem: the limiting distribution of the estimator is not centered at zero, and standard test statistics (e.g., Wald, LM, and LR) no longer follow their conventional asymptotic $\chi^{2}$ distributions. To address this, we propose an analytical bias-correction procedure that restores valid large-sample inference for both parameter estimates and test statistics.
@@
Notes: Model. $Y_{it}=\mathds{1}\{\rho Y_{it-1}+(\beta+\alpha_{1,i}+\gamma_{1,t})Z_{it}+\alpha_{2,i}+\gamma_{2,t}+\varepsilon_{it}>0\}$ where $\mathds{1}\{\cdot\}$ denotes the indicator function, $\varepsilon_{it}$ is standard logistic independent of the exogenous regressor $Z_{it}$, and the true values $\rho_{0}=\beta_{0}=0.5$. $\widehat{\beta}$ and $\widehat{\rho}$ are the uncorrected estimators while $\widehat{\beta}_{L}$ and $\widehat{\rho}_{L}$ are the corrected estimators. $1000$ replications. The detailed data-generating process is given in Section 4.
@@
As a preview, Figure 1 presents simulated boxplots of the estimation bias in a two-way heterogeneous parameter logit model, comparing estimators from our bias correction procedure to the uncorrected ones. The results reveal that our bias correction procedure significantly reduces the estimation bias while maintaining comparable mean squared errors.
@@
Literature Review. The estimation and inference of fixed-effects models in the presence of the incidental-parameter problem have been extensively studied. Early work, such as Chamberlain (1980), Arellano and Bond (1991), and Lancaster (2002), focused on frameworks with only individual effects in the intercept, establishing fixed-$T$ consistency (short panels) for structural parameter estimators in specific models. In recent decades, the availability of long panel datasets has motivated a large–$N,T$ framework, where the cross-sectional size $N$ and the time dimension $T$ grow at similar rates under rectangular asymptotics. In this setting, estimators remain consistent but exhibit a non-negligible bias of order $O\left(1/T\right)$ because each individual effect is estimated from only $T$ observations. This incidental-parameter bias is especially pronounced in nonlinear or dynamic models. To address it, researchers have developed a variety of bias correction techniques within a maximum likelihood framework. Hahn and Newey (2004) and Hahn and Kuersteiner (2011) propose “parameter-based” corrections that remove the bias from the likelihood estimator; Woutersen (2002) and Li et al. (2003) propose “score-based” methods that modify the profiled score; while “likelihood-based” approaches such as Bester and Hansen (2009) and Arellano and Hahn (2016) adjust the log-likelihood directly. Most of these procedures are analytical, relying on closed-form approximations to the bias. Alternatively, numerical corrections estimate the bias through resampling or re-estimation, including the jackknife (Dhaene and Jochmans, 2015), the bootstrap (Bellio et al., 2023; Higgins and Jochmans, 2024), and integrated-likelihood methods (Arellano and Bonhomme, 2009).
@@
For two-way fixed effects models, incorporating time effects in the intercept generates an additional bias of order $O\left(1/N\right)$ on top of the $O\left(1/T\right)$ term from individual effects. Several studies extend bias-correction methods to this setting. Moon and Weidner (2017) develop a parameter-based analytical correction for dynamic linear models with interactive effects. For nonlinear models, Fernández-Val and Weidner (2016) propose parameter-based techniques for additive fixed effects, while Chen et al. (2021) study interactive effects in static models. Alternatively, Jochmans and Otsu (2019) provide a likelihood-based approach for static models with an arbitrary but known fixed-effect structure. Other model-specific contributions include Bai (2009) and Charbonneau (2017). For a comprehensive overview of these developments, see Fernández-Val and Weidner (2018). Despite this progress, the two-way literature primarily addresses heterogeneity in levels—that is, intercept effects—rather than heterogeneity in responses (slopes). Extending bias-correction methods to settings with slope heterogeneity remains largely unexplored and is the focus of our contribution.
A substantial related literature explores heterogeneity in slope coefficients within panel models, emphasizing individual-specific parameters (Robertson and Symons, 1992; Pesaran and Smith, 1995) and time-varying coefficients (Robinson, 1989; Stock and Watson, 1996), most often in linear or static contexts (Hsiao, 2014; Pesaran, 2015). Recent research in dynamic nonlinear panels has advanced these ideas further. Chernozhukov et al. (2013) derive partial identification with uniform inference for nonseparable panels under time-homogeneity. Browning and Carro (2014) provide mixture-based point identification for dynamic binary models with maximal cross-sectional heterogeneity, while Browning and Carro (2010) analyze short-panel dynamic binary models with heterogeneous transition probabilities and propose a mean-integrated-MSE (MIMSE) estimator that better balances bias–variance trade-offs in small–$T$ settings than analytical bias correction. While these studies focus on identification or fixed–$T$ settings within semiparametric or nonparametric frameworks, we adopt a parametric, likelihood-based approach that achieves point identification and enables bias-corrected inference for dynamic nonlinear models with two-way slope and intercept effects under large–$N,T$ asymptotics.
Two recent studies are closely aligned with our framework. Keane and Neal (2020) examine linear models with two-way slope and intercept effects, employing an iterative “mean-observation OLS” estimator to address bias. Their analysis of U.S. agricultural data reveals pronounced regional differences in heat-yield sensitivity, alongside temporal adaptation as farmers adopt new technologies over time. Lu and Su (2023) study a similar two-way heterogeneous-slope specification and implement a parameter-based jackknife estimator that enables uniform inference for structural parameters. Their cross-country analysis of the Feldstein–Horioka relation uncovers variation across nations and periods, driven by financial integration and evolving policy regimes. Both studies underscore the importance of slope heterogeneity across individuals and time but remain confined to linear settings. Our study advances this literature by extending the framework to a dynamic nonlinear context and developing a likelihood-based analytical bias-correction procedure that ensures valid estimation and inference. To our knowledge, neither our DN-HP model nor its associated bias correction has been previously explored. Together, the model and method provide a unified approach to estimation and inference in dynamic panels with multi-dimensional heterogeneity.
@@
This paper. We employ the maximum likelihood framework and focus on multi-dimensional additive two-way fixed effects, whose number grows with the sample size. We consider an arbitrary DN-HP model whose log-likelihood function is specified up to the unknown parameters. We construct a modified, or corrected, log-likelihood function by adding two bias correction terms to the original one. These two bias terms are analytically derived by combining i) the Taylor expansion of the original log-likelihood function, in terms of the fixed effects, and ii) an asymptotic expansion of the fixed effects themselves. We show that the corrected log-likelihood function does not suffer from the incidental parameter problem, under appropriate regularity conditions. Our method can be viewed as an extension of Arellano and Hahn (2016) and Bester and Hansen (2009) to the two-way DN-HP models.
@@
One key benefit of our procedure is that we achieve bias corrections for the point estimators and the test statistics by modifying a single object, the log-likelihood function. We rigorously show that the estimators obtained by maximizing the corrected log-likelihood function retains a limiting normal distribution with zero mean, consistent with the asymptotic theory of the classical maximum likelihood estimators (MLE) in the absence of incidental parameters. Beyond the estimators, we also show that the likelihood ratio (LR), Lagrange-multiplier (LM), and Wald statistics, derived from the same corrected log-likelihood function, are asymptotically equivalent, sharing the same asymptotic $\chi^{2}$ distribution under the null hypothesis.
@@
We demonstrate the finite-sample performance of our method through Monte Carlo simulations, comparing it to the original likelihood and potential bias correction devices of the bootstrap and the jackknife. We show that our correction procedure reduces the bias significantly, without increasing the root mean squared errors, and restores the test sizes of the LR test to its nominal level. We find that our procedure outperforms the jackknife when it comes to heterogeneous slope coefficients. We also find that the LR test based on our approach has a stronger power near the true hypothesis than the LR test based on the bootstrap. In the empirical application, we apply our approach with a probit model to study the aforementioned question of how the number of children affects a mother’s labor force participation, with a focus on single-mother households. Our analysis reveals that this impact varies significantly across individuals and over time. Neglecting such variation can lead to biased estimates and markedly different conclusions, highlighting the need to address both level and response heterogeneities, as well as the value of our framework in empirical research.
@@
The remainder of the paper is organized as follows. Section 2 explains our settings, the identification restriction, and the incidental parameter problem of DN-HP models. Section 3 describes our bias correction procedure and provides relevant statistical properties. Section 4 presents simulation studies to demonstrate the performance of the method. Section 5 applies our method to study single mothers’ labor force participation. Finally in Section 6, we leave some closing remarks. Proofs, technical details, additional results, elaborated discussions, etc. are provided in the appendix. These items have references starting with letters.
@@ 2.1 Models and Estimations
$Y_{it}|X_{it},\ldots,X_{i1},\phi^{0}\sim f(y|X_{it},\theta_{0}+\alpha_{i}^{0}+\gamma_{t}^{0}),$ (2.1)